OpenStudy (anonymous):

If 6x-3y=30 and 4x=2-y then find x+y

OpenStudy (zzr0ck3r):

x =( 30 - 3y)/6 y = 2 - 4x x + y = (30-3y)/6 + (2-4x)

OpenStudy (anonymous):

x=(30-3(2-4x))/6 x=(30-6+12x)/6 x=(24+12x)/6 x=4+2x -x=4 so x=-4 and y=2-4x so y=2-4*(-4) y=2+16 y=18

OpenStudy (anonymous):

x+y= -4+18 x+y=14

OpenStudy (anonymous):

\[\{6 x-3 y=30,4 x=2-y\}\]The solutions to the above equations is: x=2 and y=-6 x+y=-4 - - - - - - - - - - - - 4 x = 2 - y y = 2 - 4x 6 x - 3 y = 30 6x - 3(2 - 4x) = 30 6x - 6 + 12x = 30 18x = 36 x = 2 y = 2 - 4x y = 2 - 4*2 y = -6